# Zero-Shot Temporal Resolution Domain Adaptation for Spiking Neural Networks

### v2

@@ Abstract
Spiking Neural Networks (SNNs) are biologically-inspired deep neural networks that efficiently extract temporal information while offering promising gains in terms of energy efficiency and latency when deployed on neuromorphic devices. However, SNN model parameters are sensitive to temporal resolution, leading to significant performance drops when the temporal resolution of target data during deployment is not the same with that of the pre-deployment source data used for training, especially when fine-tuning with new data is not possible during deployment. To address this challenge, we propose three novel domain adaptation methods for adapting neuron parameters to account for the change in time resolution without re-training on target time-resolution. The proposed methods are based on a mapping between neuron dynamics in SNNs and State Space Models (SSMs) and are applicable to general neuron models. We evaluate the proposed methods under spatio-temporal data tasks, namely the audio keyword spotting datasets SHD and MSWC as well as the neuromorphic image classification NMINST dataset. Our methods provide an alternative to - and in majority of the cases significantly outperform - the existing reference method that simply scales the time constant. Moreover, our results show that high accuracy on high temporal resolution data can be obtained by time efficient training on lower temporal resolution data and model adaptation.
@@ 1 Introduction
Domain adaptation addresses the scenarios where a model is confronted by a domain shift after deployment, while its task remains the same, such as face recognition under different lightning conditions or poses [domain_adaptation_brief_review, liang2023comprehensivesurveytesttimeadaptation]. The domain shift refers to a change in feature space between the source domain on which the model was trained, and the target domain on which the model is evaluated. The change in feature space has been studied from various aspects, particularly for visual data [Patel_2015, zubić2024state], but has rarely been explored with respect to temporal resolution shift for temporal data. Our work focuses precisely on this under-explored aspect, investigating feature discrepancies caused by differences in temporal resolution between the source and target data. While domain adaptation typically involves transferring knowledge between the source and target data by re-training parts of the model or the entire model using samples from target domain, [Patel_2015, Sun_2019_CVPR], we focus on the extreme case of no retraining on target domain.
In a conventional data acquisition set-up, where values of signals of interests are recorded at regular intervals, temporal resolution is primarily determined by the time interval between successive samples. For instance, the frame rate of a video stream or the sampling rate of an audio recording determines the temporal resolution. In the case of event-based sensing, e.g. [Gallego2022], where changes in the scene are recorded with time-stamps, temporal resolution is determined by the temporal accuracy of time-stamps.
Various practical limitations lead to data with varying temporal resolution. For instance, limited memory capacity and power constraints often require lower sampling rates of the data acquisition or storage devices [Dieter_2005, ur2016BigDataCompress]. Additionally, the quality of sensors and bandwidth limitations further restrict temporal resolution, as high-resolution sensors and real-time transmission demand more resources [chen2021distributed, park2021communication]. Consequently, changes in the temporal resolution of data has been the focus of a wide range of studies, such as increasing of temporal resolution of videos [Zuckerman2020], optimization of sensor networks’ sensing rates [Ankur2004], processing of irregularly sampled data [rubanova2019latent], optimal combination of high- and low-rate time-series data [Bae2024]. In the case of event-based cameras, effect of bandwidth and event-rate constraints on the temporal resolution have been studied [gehrig2022highresolutioneventcamerasreally, zubić2024state].
@@
Figure 1: Overview of the set-up. We investigate how models trained on data with a given source temporal resolution can be adapted for data with a different target temporal resolution, both in the Fine-to-Coarse and Coarse-to-Fine deployment directions. The diagram illustrates Coarse-to-Fine scenario.
@@
Applications of these scenarios appear at different edge processing scenarios, where small resource-constrained devices with different sensing and computing capabilities must process their data at the edge of a wireless system with varying degrees of support from a central processing or coordination unit [chen2021distributed, park2021communication]. A possible application of Coarse-to-Fine deployment is encountered in communication constrained distributed scenarios with a central learning unit. In such scenarios, sensor data is sent to a central learning unit with reduced temporal resolution due to bandwidth or power limitations [chen2021distributed, park2021communication]. As a result, the central model is trained on coarse data and then deployed to edge devices, where higher-resolution data is available for inference. A possible application of Fine-to-Coarse deployment is encountered when the sensors have to switch to low sampling rates to reduce power consumption [Dieter_2005, Ankur2004]. This leads to having target data at edge with coarser temporal resolution than the source data used for training the deployed models at the central learning unit.
@@
We consider the above temporal domain adaptation problems under Spiking Neural Networks (SNN) models. SNNs are biologically-inspired neural networks with stateful recurrent neurons and binary spike-based communication [MAASS19971659]. SNNs provide an attractive framework for processing data with encoded temporal information due to their ability to represent and learn spatiotemporal features [bittar2022surrogate, yik2024neurobench, Braille_SNN, eshraghian2023trainingspikingneuralnetworks, ma2025spikingneuralnetworkstemporal]. Like most machine learning algorithms, the current mainstream pipeline for SNNs relies on training on high-power computational units, mostly GPUs, and then deploying and running on low-power edge devices [10177729]. SNNs, when deployed at neuromorphic edge hardware, are especially valuable because they are designed to mimic the brain’s energy-efficient processes, offering significant reductions in power consumption and latency [9395703, caccavella2023lowpower, Braille_SNN]. SNNs are particularly effective in event-based systems, where event-based sensors, e.g. event-cameras record changes only as they occur, rather than at fixed time intervals, leading to optimized energy use and reduced data redundancy [Gallego2022].
@@
A common practical training set-up for SNNs is the standard digital clock-based SNNs, where the temporal resolution of the data corresponds directly to the processing resolution at which the model operates. In this setup, the possible discrepancies in the temporal resolution of the source and target data constitutes a significant challenge since trained SNN model parameters are strongly dependent on the temporal resolution of source data. Hence, an SNN model that has been trained using data with a given temporal resolution may not necessarily perform well in inference on data from another sensory source even if it is recording of the same phenomena, e.g., see Section LABEL:sec:numericalResults.
@@
Training on sequences with high temporal resolution is desired due to possible inclusion of more information and hence better performance of the model [gehrig2024low]. However this comes with a cost as difficulties arise when training SNNs on long sequences such as exploding and vanishing gradients and other instabilities due to the recurrent and saturating nature of SNNs [zucchet2024recurrentneuralnetworksvanishing]. Another limitation while training on long sequences is the significant training time as the temporal resolution dimension linearly increases the training time and computational complexity [pereznieves2022sparse]. Hence, the development of temporal adaptation methods enables training models in a fast and stable manner on low temporal resolution data, while still ensuring high performance during inference on high temporal resolution data. This setting supports energy, time and computation efficient training pipeline.
@@
Considering both the temporal domain adaptation problem as well as challenges of training on sequences with fine temporal resolution, the main contributions of this paper are as follows:
@@
  • The performance of the proposed methods is demonstrated using the audio datasets SHD [shddataset] and MSWC [mswcdataset] and the vision NMNIST dataset [orchard2015converting] .
@@
  • We show that utilizing low temporal resolution version of the data for pre-training constitutes a promising training approach for improving computational training efficiency in SNNs.
@@ 2 Related Work
The Leaky-Integrate-and-Fire (LIF) neuron model is one of the most simple and widely used neuron models in the context of SNNs. When input stimuli in form of spikes are applied to the neuron, the membrane potential increases and when a certain threshold is reached it produces an output spike. On the other hand, in the absence of input stimuli, the membrane potential decays exponentially [bittar2022surrogate], see Figure 2. The vanilla LIF neuron has only one learnable parameter, which is the membrane potential decay (the leak), and this parameter has explicit exponential dependence on the time step length of the data. Hence when dealing with change of temporal resolution, intuitive way to adapt the LIF neuron dynamic is to scale the leak parameter according to the change. In [he2020comparing] the authors explore this method of temporal domain adaptation using the simple LIF neuron. On the other hand, in [caccavella2023lowpower], the authors use Integrate-and-Fire (IF) neuron which is a special simpler case of LIF neuron without leak. Hence alternative to parameters adaptation, they employ a graded spike network and different normalization techniques for domain adaptation with promising results. However, in order to capture more complex temporal information in data, more complex neuron models such as the adaptive LIF (adLIF) neurons have shown superior performance [bittar2022surrogate, baronig2025advancingspatiotemporalprocessingspiking, adLIFNature, ma2025spikingneuralnetworkstemporal]. In the case of more complex neuron models however, the parameter might not have explicit dependence on the time-length of the data, or even any dependence at all.
@@
An important building block of our proposed approach is State Space Models (SSMs), which is a general framework for modelling dynamic systems. Their neural network implementations, i.e. deep SSMs, raised as an attractive alternative to the conventional deep recurrent models. It was shown that for appropriate conditioning on the state matrix dynamics, deep SSMs can handle long-range dependencies while offering stable and computationally efficient training [wang2024state]. Limited explorations of temporal resolution change have been performed using SSMs and traditional data [smith2022simplified, Table 2]. Deep SSMs have been used for event-based data in a number of works [schöne2024scalable, zubić2024state] Similar to the leak adaptation in LIF neurons, explicit correlation can be addressed between significantly simplified version of SSM and the temporal dynamics between spikes. [schöne2024scalable] explores this concept by processing incoming spikes with adapted leak based on the time-length between consecutive spikes. Alternatively, [zubić2024state] explored varying-resolution aware training of deep SSMs for vision dataset, where the training includes masking inputs outside specific frequency ranges and incorporating an additional loss function to penalize high-frequency activities.
Interesting intersection of work is emerging between SNNs and deep SSMs. One common line of work is embedding spiking dynamics in deep SSMs where at each time step the output activity of the SSM is processed by a firing threshold function [stan2023learning, bal2024rethinkingspikingneuralnetworks, karilanova2025statespacemodelinspiredmultipleinput]. Alternatively, in [du2024spiking] they combine SNN with SSMs by adding a spiking layer to the SSM. Nevertheless, the challange of drawing lines of correspondence between SNNs and linear SSMs is still present due to the non-linearity when it comes to SNNs. Our proposed methods for temporal domain adaptation for SNNs are based on this drawn parallel and the theory developed for adaptation in SSMs.
@@ 3.1 Popular spiking neuron model
The discrete-time adaptive Leaky-Integrate and Fire (adLIF) neuron is defined as [bittar2022surrogate]
@@
where $\bm{s}_{in}\in\mathbb{Z}^{L_{b}\times 1}$ are the input spikes from the neurons in the previous layer and $\bm{W}\in\mathbb{R}^{1\times L_{b}}$ is the weight matrix associated with their connections to the neuron of observation, while $\bm{s}_{out}\in\mathbb{Z}^{L_{c}\times 1}$ are the output spikes of all the other neurons in the current layer and $\bm{V}\in\mathbb{R}^{1\times L_{c}}$ with zero diagonal is their connections to the neuron of observation. Here, $L_{c}$ and $L_{b}$ are the number of neurons in the current and the previous layer, respectively. Furthermore, $g_{\theta}$ describes the spiking behavior with firing threshold $\theta\in\mathbb{R}^{1\times 1}$; $\alpha=\exp(\Delta/\tau_{mem})$ represents the decay of the membrane potential over time such that $\Delta$ corresponds to the duration of a timestep and $\tau_{mem}$ is a membrane time constant; $\beta=\exp(\Delta/\tau_{syn})$ represents the decay of the adaptive variable with adaptation constant $\tau_{syn}$; the parameters $a$ and $b$ are without explicit dependence on the discrete time step length $\Delta$. All parameters $\alpha,\beta,a,b$ are generally learnable during training within given ranges.
Note, the weight vector $\bm{V}$ corresponds to recurrent connections between the neurons in the same layer. Recurrence represented by $\bm{V}$ is a characteristic of the network rather than the neuron.
@@ 3.2 General spiking neuron model
In this section we introduce a discrete-time general n-dimensional spiking neuron model which generalizes the two-dimensional adLIF neuron (1). This generalized model is central to the proposed methods of this paper. It is defined as
@@
where $\bm{v}\in\mathbb{R}^{n\times 1}$ is the state variable, while $\bm{H}_{v}\in\mathbb{R}^{n\times n}$ and $\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\in\mathbb{R}^{n\times 1}$ are matrices containing the parameters of the individual neuron. The subscripts of the $\bm{H}$ matrices, i.e. $v,f,i$, and $r$, indicate state, feedback, input and recurrence, respectively. The spiking behaviour is described by $g_{\bf{\Theta}}(\cdot)$, where the neuron spikes when the state enters into the region described by the region $\bf{\Theta}$.
We note that the adLIF neuron can be expressed as
@@
Thus, the adLIF neuron (1) with two state variables $\bm{v}=\begin{bmatrix}u,v\end{bmatrix}^{T}$ is a special case of the general model where
@@
$\bm{H}_{r}$ $=\begin{bmatrix}1-\alpha\\ 0\end{bmatrix}.$

### v3

@@ Abstract
Spiking Neural Networks (SNNs) are biologically-inspired deep neural networks that efficiently extract temporal information while offering promising gains in terms of energy efficiency and latency when deployed on neuromorphic devices. SNN parameters are sensitive to temporal resolution, leading to significant performance drops when the temporal resolution of target data during deployment is not the same as that of the source data used for training, especially when fine-tuning with the target data is not possible during deployment. To address this challenge, we propose three novel domain adaptation methods for adapting neuron parameters to account for the change in time resolution without re-training on target time resolution. The proposed methods are based on a mapping between neuron dynamics in SNNs and State Space Models (SSMs) and are applicable to general neuron models. We evaluate the proposed methods under spatio-temporal data tasks, namely the audio keyword spotting datasets SHD and MSWC, and the neuromorphic image NMINST dataset. Our methods provide an alternative to-and in most cases significantly outperform-the existing reference method that consists of scaling only the time constant. Notably, when the temporal resolution of the target data is double that of the source data, applying one of our proposed methods instead of the benchmark achieves classification accuracy of $89.5\%$ instead of $53.0\%$ on SHD, $93.6\%$ instead of $38.8\%$ on MSWC and $98.5\%$ instead of $97.2\%$ on NMNIST. Moreover, our results show that high accuracy on high temporal resolution data can be obtained by time-efficient training on lower temporal resolution data.
@@ 1 Introduction
Figure 1: Overview of the set-up. We investigate how models trained on data with a given source temporal resolution can be adapted for data with a different target temporal resolution, both in the Fine-to-Coarse and Coarse-to-Fine deployment directions.
Domain adaptation addresses the scenarios in which a model is confronted with a domain shift after deployment, while its task remains the same, such as face recognition under different lightning conditions or poses [18, 36]. The domain shift refers to a change in feature space between the source domain on which the model was trained, and the target domain on which the model is evaluated. The change in feature space has been studied from various aspects, particularly for visual data [47, zubić2024state], but has rarely been explored with respect to temporal resolution shift for temporal data. Our work focuses precisely on this under-explored aspect, investigating feature discrepancies caused by differences in temporal resolution between the source and target data. While domain adaptation typically involves transferring knowledge between the source and target data by re-training parts of the model or the entire model using samples from target domain, [47, 54] [3], we focus on the extreme case of no retraining on target domain.
In a conventional data acquisition set-up, where values of signals of interest are recorded at regular intervals, temporal resolution is primarily determined by the time interval between successive samples. For instance, the frame rate of a video stream or the sampling rate of an audio recording determines the temporal resolution. In the case of event-based sensing, e.g. [20], where changes in the scene are recorded with time-stamps, temporal resolution is determined by the temporal accuracy of time-stamps.
Various practical limitations lead to data with varying temporal resolution. For instance, limited memory capacity and power constraints often require lower sampling rates of data acquisition or storage devices [15, 55]. Additionally, sensor quality and bandwidth limitations further restrict temporal resolution, as high-resolution sensors and real-time transmission demand more resources [10, 46]. Consequently, changes in the temporal resolution of data have been the focus of a wide range of studies, such as increasing of temporal resolution of videos [64], optimization of sensor networks’ sensing rates [31], processing of irregularly sampled data [49], optimal combination of high- and low-rate time-series data [2]. In the case of event-based cameras, effect of bandwidth and event-rate constraints on the temporal resolution have been studied [21, zubić2024state].
@@
Applications of these scenarios appear in different edge processing scenarios, where small resource-constrained devices with different sensing and computing capabilities must process their data at the edge of a wireless system with varying degrees of support from a central processing or coordination unit [10, 46]. A possible application of Coarse-to-Fine deployment is encountered in communication constrained distributed scenarios with a central learning unit. In such scenarios, sensor data is sent to a central learning unit with reduced temporal resolution due to bandwidth or power limitations [10, 46]. As a result, the central model is trained on coarse data and then deployed to edge devices, where higher-resolution data is available for inference. A possible application of Fine-to-Coarse deployment is encountered where the sensors have to switch to low sampling rates to reduce power consumption [15, 31]. This leads to having target data at edge with coarser temporal resolution than the source data used for training the deployed models at the central learning unit.
@@
We consider the above temporal domain adaptation problems under Spiking Neural Networks (SNN) models. SNNs are biologically-inspired neural networks with stateful recurrent neurons, and binary spike-based communication [39]. SNNs provide an attractive framework for processing data with encoded temporal information due to their ability to represent and learn spatiotemporal features [6, 60, 41, 17, 38]. Like most machine learning algorithms, the current mainstream pipeline for SNNs relies on training at high-power computational units, such as GPUs, and then deploying and running on low-power edge devices [1]. SNNs, when deployed at neuromorphic edge hardware, are especially valuable because they are designed to mimic the brain’s energy-efficient processes, offering significant reductions in power consumption and latency [14, 7, 41]. SNNs are particularly effective in event-based systems, where event-based sensors, e.g. event-cameras, record changes only as they occur, rather than at fixed time intervals, leading to optimized energy use and reduced data redundancy [20].
@@
A common practical training set-up for SNNs is the standard digital clock-based SNNs, where the temporal resolution of the data corresponds directly to the processing resolution at which the model operates. In this setup, the possible discrepancies in the temporal resolution of the source and target data constitute a significant challenge since trained SNN model parameters are strongly dependent on the temporal resolution of the source data. Hence, an SNN model that has been trained using data with a given temporal resolution may not necessarily perform well in inference on data from another sensory source even if it is recording of the same phenomena, e.g., see Section 5.
@@
Training on sequences with high temporal resolution is desired due to the possible inclusion of more information and therefore better performance of the model [22]. However, this comes with a cost. Difficulties arise when training SNNs on long sequences such as exploding and vanishing gradients and other instabilities due to the recurrent and saturating nature of SNNs [63]. Another limitation while training on long sequences is the significant training time as the temporal resolution dimension approximately linearly increases the training time and computational complexity [48]. Hence, the development of temporal adaptation methods enables training models in a fast and stable manner on low temporal resolution data, while still ensuring high performance during inference on high temporal resolution data. This setting supports energy, time, and computation efficient training pipeline.
@@
Considering both the temporal domain adaptation problem as well as challenges of training on sequences with fine temporal resolution, the main contributions of this paper are:
@@
  • The performance of the proposed methods is demonstrated using the audio datasets SHD [12] and MSWC [40] and the vision NMNIST dataset [45] .
@@
  • We show that utilizing version of the data with low temporal resolution for pre-training, constitutes a promising training approach for improving computational training efficiency in SNNs.
We now provide an overview of the organization of the paper. Section 2 provides a review of the related work. The main problem studied in this article, i.e., the temporal domain adaptation problem, is formulated in Section 3. Section 4 presents the proposed methods. Numerical results illustrating the performance of our methods are provided in Section 5 and Section 6. Finally, in Section 7 and Section 8 we present further discussions and a summary of our conclusions, respectively.
@@ 2 Related Work
The Leaky-Integrate-and-Fire (LIF) neuron model is one of the most simple and widely used neuron models in the context of SNNs. When input stimuli in the form of spikes are applied to the neuron, the membrane potential increases and when a certain threshold is reached it produces an output spike. On the other hand, in the absence of input stimuli, the membrane potential decays exponentially [6], see Figure 2. The vanilla LIF neuron has only one learnable parameter, which is the membrane potential decay (the leak), and this parameter has an explicit exponential dependence on the time step length of the data. Hence, when dealing with a change of temporal resolution, intuitive way to adapt the LIF neuron dynamic is to scale the leak parameter according to the change. In [25, Table 12] the authors explore this method of temporal domain adaptation using the simple LIF neuron. On the other hand, in [7], the authors use Integrate-and-Fire (IF) neuron which is a special simpler case of LIF neuron without leak. Hence alternative to parameters adaptation, they employ a graded spike network and different normalization techniques for domain adaptation with promising results. However, in order to capture more complex temporal information in data, more complex neuron models such as the adaptive LIF (adLIF) neurons have shown superior performance [6, 5, 61, 38]. In the case of more complex neuron models however, the parameter might not have explicit dependence on the time-length of the data, or even any dependence at all.
@@
An important building block of our proposed approach is State Space Models (SSMs), which is a general framework for modeling dynamic systems. Their neural network implementations, i.e. deep SSMs, raised as an attractive alternative to the conventional deep recurrent models. It was shown that for appropriate conditioning on the state matrix dynamics, deep SSMs can handle long-range dependencies while offering stable and computationally efficient training [57]. Limited explorations of temporal resolution change have been performed using SSMs and traditional data [52, Table 2]. Deep SSMs have been used for event-based data in a number of works [schöne2024scalable, zubić2024state] Similar to the leak adaptation in LIF neurons, explicit correlation can be addressed between significantly simplified version of SSM and the temporal dynamics between spikes. [schöne2024scalable] explores this concept by processing incoming spikes with adapted leak based on the time-length between consecutive spikes. Alternatively, [zubić2024state] explored varying-resolution aware training of deep SSMs for vision dataset, where the training includes masking inputs outside specific frequency ranges and incorporating an additional loss function to penalize high-frequency activities.
Interesting intersection of work is emerging between SNNs and deep SSMs. One common line of work is embedding spiking dynamics in deep SSMs where at each time step the output activity of the SSM is processed by a firing threshold function [56, 4, 33, 32]. Alternatively, in [16] they combine SNN with SSMs by adding a spiking layer to the SSM. Nevertheless, the challenge of drawing lines of correspondence between SNNs and linear SSMs is still present due to the non-linearity when it comes to SNNs. Our proposed methods for temporal domain adaptation for SNNs are based on this drawn parallel and the theory developed for adaptation in SSMs. See Table 1 for a summary of the prior works discussed in this section.
Method Neuron model Adaptation Datasets Metrics
This article Generalized spiking neuron including adLIF, LIF, IF, SSM Expectation, Integral, and Euler Methods SHD, MSWC, N-MNIST Classification accuracy in varying temoral resolution
Comparing SNNs and ANNs [25, Table 12] LIF, rate-coded RNN and rate-coded LSTM No-adaptation N-MNIST Classification accuracy in varying temporal resolution
Event-driven inference [7] IF Layer normalization class “faces” of N-CALTECH101 mAP on varying time-window of integration
EventSSM [schöne2024scalable] SSM Event-by-event processing SHD, SSC, DVS-Gesture Classification accuracy (both in train and test stage event-by-event processing used)
SSM for event cameras [zubić2024state] SSM Low-pass bandlimiting, and SSM-ViT block Gen1 and 1 Mpx mAP on varying evaluations frequency
Table 1: Overview of works which include aspects of zero-shot temporal adaptation in SNN or deep-SSM models.
@@ 3.1 Popular spiking neuron model
The discrete-time adaptive Leaky-Integrate and Fire (adLIF) neuron is defined as [6]
@@
where $\bm{s}_{in}\in\mathbb{Z}^{L_{b}\times 1}$ are the input spikes from the neurons in the previous layer and $\bm{W}\in\mathbb{R}^{1\times L_{b}}$ is the weight matrix associated with their connections to the neuron of observation, while $\bm{s}_{out}\in\mathbb{Z}^{L_{c}\times 1}$ are the output spikes of all the other neurons in the current layer and $\bm{V}\in\mathbb{R}^{1\times L_{c}}$ with zero diagonal is their connections to the neuron of observation. Here, $L_{c}$ and $L_{b}$ are the number of neurons in the current and the previous layer, respectively. Furthermore, $g_{\theta}(\cdot)$ describes the spiking behavior with the firing threshold $\theta\in\mathbb{R}^{1\times 1}$; $\alpha=\exp(\Delta/\tau_{mem})$ represents the decay of the membrane potential over time such that $\Delta$ corresponds to the duration of a timestep and $\tau_{mem}$ is a membrane time constant; $\beta=\exp(\Delta/\tau_{syn})$ represents the decay of the adaptive variable with adaptation constant $\tau_{syn}$; the parameters $a$ and $b$ are without explicit dependence on the discrete time step length $\Delta$. All parameters $\alpha,\beta,a,b$ are generally learnable during training within given ranges.
Note, the weight vector $\bm{V}$ corresponds to the recurrent connections between the neurons in the same layer. Recurrence represented by $\bm{V}$ is a characteristic of the network rather than the neuron.
@@ 3.2 General spiking neuron model
In this section, we introduce a discrete-time general $n$-dimensional spiking neuron model, which generalizes the two-dimensional adLIF neuron (1). This generalized model is central to the proposed methods of this paper. It is defined as
@@
where $\bm{v}\in\mathbb{R}^{n\times 1}$ is the state variable, while $\bm{H}_{v}\in\mathbb{R}^{n\times n}$ and $\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\in\mathbb{R}^{n\times 1}$ are matrices containing the parameters of the individual neuron. The subscripts of the $\bm{H}$ matrices, i.e. $v,f,i$, and $r$, indicate state, feedback, input and recurrence, respectively. The spiking behavior is described by $g_{\bf{\Theta}}(\cdot)$, where the neuron spikes when the state enters in the region described by $\bf{\Theta}$. The region $\bf{\Theta}$ depends on the neuron model.
Many popular spiking neuron models can be expressed as special cases of the general spiking neuron model in (2) including the LIF neuron [23] with $n=1$ state variable, the adLIF neuron [6] with $n=2$, the Izikevich neuron [30] with $n=2$, the Hodgkin-Huxley [28] with $n=4$, multi-compartment neurons with $2$ compartments each with $n=2$ [62]. Below, we show how the $n=2$ dimensional adLIF neuron can be expressed as a special case of (2).
The adLIF neuron, in matrix form, can be written as:
@@
Thus, the adLIF neuron (1) with two state variables $\bm{v}=\begin{bmatrix}u,v\end{bmatrix}^{T}$ is a special case of the general model in (2) where
@@
$\bm{H}_{r}$ $=\begin{bmatrix}1-\alpha\\ 0\end{bmatrix}$. (4)
Remark 1.
The general spiking neuron model in (2) does not encompass all possible neuron formulations. For instance, it does not capture: (i) hard reset mechanism, where the neuron state is set to a predefined reset state value when spiking occurs [13] rather then subtracting the spike using $\bm{H}_{f}$; (ii) varying adaptive spiking threshold, where the threshold varies over time [11], rather than being defined by a fixed spiking region $\bf{\Theta}$.